# Public Projects with Preferences and Predictions

### v1

@@ 2 Quadratic Transfers Mechanism
In this section, we give results, including a Price of Anarchy bound, for the Quadratic Transfers Mechanism (QTM) for the public projects problem. These results complement existing analyses [Eguia et al., 2019, 2023], and will also serve as a foundation for our results in later sections combining predictions and preferences.
@@ 2.1 Model
Public projects.
We consider a classic public projects setting in which a group of $n$ agents needs to select one of $m\geq 2$ alternatives. We will use $i,j$ to denote generic agents and $k,l$ to denote generic alternatives. Let $v_{k}^{i}\in\operatorname{\mathbb{R}}_{\geq 0}$ represent agent $i$’s value for alternative $k$. We assume that all values are nonnegative. Denote $\mathbf{v}^{i}=(v_{1}^{i},v_{2}^{i},\ldots,v_{m}^{i})$ as the vector of agent $i$’s values across all $m$ alternatives. The aggregate value for alternative $k$ is $V_{k}:=\sum_{i=1}^{n}v_{k}^{i}$. Without loss of generality, we number the alternatives such that $V_{1}\geq V_{2}\geq\ldots\geq V_{m}$.
Mechanisms and welfare.
A mechanism defines a game played by the set of agents. Given the actions of the agents, the output of the mechanism is a probability distribution $\mathbf{p}\in\Delta_{m}$, the set of distributions over the alternatives, along with a net payment from each agent. We assume quasilinear utility: If the mechanism’s output is $\mathbf{p}$ and if agent $i$ makes a net payment $\pi_{i}$, then $i$’s expected utility is given by $\sum_{k}p_{k}v_{k}^{i}-\pi_{i}$. The social welfare of the mechanism is $\sum_{k}p_{k}V_{k}$, the expected aggregate value produced. The optimal or “first-best” social welfare is $V_{1}$, the aggregate value of the best alternative.
@@
In the Quadratic Transfers Mechanism (QTM) [Eguia et al., 2019, 2023], each agent $i$ simultaneously and privately submits a vector $\mathbf{a}^{i}\in\operatorname{\mathbb{R}}^{m}$, where the $k$th entry $a_{k}^{i}$ represents a number of “votes” for alternative $k$. Votes are allowed to be negative, representing voting against that alternative. We let $A_{k}:=\sum_{i}a_{k}^{i}$ denote the aggregate votes for alternative $k$.
@@
2.2 Equilibrium Characterization
In this subsection, we characterize conditions under which a pure-strategy equilibrium exists in the QTM. Our results build on some techniques of  Eguia et al. [2019], but also introduce new techniques and new structural features of the mechanism, such as the fact that votes across alternatives always sum to zero in equilibrium.
To show existence, we take the following steps. First (Lemma 7), we show that for each player, it is strictly dominated to play any $\mathbf{a}^{i}$ lying outside a certain compact set. It follows that all equilibria are “as if” the game were restricted to a compact strategy space (Lemma 8). From there, we can show that if the parameter $c$ is large enough relative to the maximum agent preference, then utilities are concave, and apply an existence theorem (Lemma 1). A similar argument appears in Eguia et al. [2019], but in our case we give a constructive bound on $c$ that appears quite tight. We also note a connection in the analysis to exponential family distributions. Finally, we derive the first-order conditions for pure-strategy Nash equilibrium, and show as a byproduct that each agent’s votes sum to zero.
All proofs appear in Appendix A.1.
Lemma 1.
@@ 2.2 Equilibrium Characterization
We now state the important first-order conditions that characterize equilibrium in terms of the mechanism’s expected individual and aggregate welfare, $\operatorname{\mathbb{E}}_{k\sim p}v_{k}^{i}=\sum_{k}p_{k}v_{k}^{i}$ and $\operatorname{\mathbb{E}}_{k\sim p}V_{k}=\sum_{k}p_{k}V_{k}$ respectively. This result also appears throughout Eguia et al. [2019], although we add some observations that are useful to us later in constructing equilibria as part of our mechanisms.
Lemma 2.
$A_{k}$ $=\frac{p_{k}}{2c}\left(V_{k}-\sum_{\ell}p_{\ell}V_{\ell}\right).$ (5)
Furthermore, if the mechanism sets $c\geq\tfrac{1}{2}\max_{i,k}v_{k}^{i}$, then $\mathbf{a}$ is a pure-strategy Nash equilibrium if and only if (4) and (5) are satisfied. Furthermore, for any vote totals $\mathbf{V}$, if $\mathbf{A}$ is any aggregate vote vector that satisfies (5), then for any $\mathbf{v}$ summing to $\mathbf{V}$ there exists a corresponding $\mathbf{a}$ summing to $\mathbf{A}$ satisfying (4).
An equilibrium is unique if Equation 5 has a unique solution $\vec{A}$. We do not know conditions under which there is a unique pure-strategy equilibrium, but conjecture that it is often unique. Summing over Equations 4 and 5, we obtain:
Corollary 1.
2.3 Price of Anarchy: Two Alternatives
In this section, we present our results for the QTM on $m=2$ alternatives. Conceptually similar results have already been shown by Eguia et al. [2019, 2023] (see Section 1.2), but they focus more on the challenges of the $m\geq 3$ setting and have different distributional and model assumptions than we do here. Recall that $V_{1}$ is the aggregate value of the better alternative, i.e. optimal welfare, and $V_{2}$ is the aggregate value of the other. Our main result, Theorem 1, is that the QTM’s pure-strategy Price of Anarchy (pPoA) is always at least $\tfrac{1}{2}$ and approaches $1$ rapidly as the spread of the game — the ratio of $V_{1}$ to the largest individual value — grows. We also show in Corollary 2 a faster rate of convergence to $1$ when the gap — the difference between $V_{1}$ and $V_{2}$ — grows. Finally, we will characterize the total payments of the QTM in terms of the amount of “disagreement” of the participants.
@@ 2.3 Price of Anarchy: Two Alternatives
[…]
@@
Total payments.
Proposition 2.
In the QTM with two alternatives and type profile $\mathbf{v}$, with $V_{1}>V_{2}$, define the disagreement to be $D:=\frac{\sum_{i}(v_{1}^{i}-v_{2}^{i})^{2}}{(V_{1}-V_{2})^{2}}$. Then in any pure-strategy equilibrium, [^2] Here Θ()normal-$\Theta(\cdot)$ is with respect to $V_{1}-V_{2}$ growing, or at least bounded below. The proof includes fully explicit nonasymptotic bounds.
[…]
  • For $n_{1}-n_{2}\ll\sqrt{n}$, $D=\omega(1)$ and revenue grows more rapidly.
2.4 $m>2$ Alternatives
We briefly touch on the case where there are more than two alternatives. Eguia et al. [2019, 2023] give asymptotic results showing that in their settings, with a fixed distribution of bounded agents, as $n\to\infty$ the welfare of the QTM tends to optimality. Here we can give a weaker nonasymptotic bound. The proof is not immediate and involves dividing the alternatives into a group with “high” welfare (higher than the mechanism’s expectation) and the remainder, an idea from Eguia et al. [2023] used in a different way.
Proposition 3.
[…]
It is not clear if this bound can be improved non-asymptotically. The analysis of the many-alternatives case is very complex (see Eguia et al. [2019, 2023]). As far as we currently know, it may be possible to have instances of a large “spread” $\frac{V_{1}}{\max_{i,k}v_{k}^{i}}$, perhaps exponentially large in $m$, with pure-strategy Price of Anarchy bounded away from one, or perhaps as small as $O(1/m)$.
3 Incorporating External Welfare
In this section, we propose mechanisms for public projects with external welfare, meaning estimates or projections of the impact of the project additional to the values of the agents participating in the mechanism. First, we will suppose that the external welfare impacts are fixed and known to the mechanism. In Section 4, we will use this section’s mechanism for our main problem, the case where external welfare impacts are forecasts from (manipulable) mechanisms.
3.1 Model and Mechanism
We define the problem of Public Projects with External Welfare, extending the public projects setting defined in Section 2. In addition to the agent values $\mathbf{v}=(\mathbf{v}^{1},\dots,\mathbf{v}^{n})$, for each $k\in\mathbb{Y}$ we now let $B_{k}\in\operatorname{\mathbb{R}}_{\geq 0}$ represent the impact to external welfare conditional on the $k$th alternative being chosen.
Given $\mathbf{B}\coloneqq(B_{1},\dots,B_{k})$, our mechanism will run a modified QTM to choose an alternative. Our goal is to choose an alternative which maximizes total welfare, represented by the sum $V_{k}+B_{k}$. Thus let $W_{k}\coloneqq V_{k}+B_{k}$, and assume without loss of generality that alternatives are numbered according to $W_{1}\geq W_{2}\geq\ldots\geq W_{m}$. Therefore, the optimal achievable welfare is $W_{1}$ and the welfare of a mechanism with output $\mathbf{p}$ is $\sum_{k}p_{k}W_{k}=\sum_{k}p_{k}(V_{k}+B_{k})$. We let $\text{pPoA}(\mathbf{v},\mathbf{B})$ be the pure-strategy Price of Anarchy of our mechanism on profile $(\mathbf{v},\mathbf{B})$.
Mechanism.
We propose the Synthetic Players Quadratic Transfer Mechanism for public projects with external welfare. The mechanism is inspired by the approach of creating a set of synthetic players whose preferences reflect the external welfare, then playing on their behalf in the QTM alongside the real participants. However, the mechanism itself is more concrete and based on the analysis of equilibrium conditions for the QTM.
We define two versions of the mechanism. The first is a practical variant that, unfortunately, we are unable to analyze. The second is a less-practical variant that assumes the mechanism has significant knowledge about the agent values. We give Price of Anarchy results for the second variant and conjecture that similar bounds hold for the first.
The mechanisms use the following equations adapted from the first-order conditions (Lemma 2), where $p_{k}=e^{A_{k}}/\sum_{\ell}e^{A_{\ell}}$.
$A_{k}$ $=p_{k}\left(W_{k}-\sum_{\ell}p_{\ell}W_{\ell}\right)$ $(\forall k),$ (6)
$a_{k}^{\text{mech}}$ $=p_{k}\left(B_{k}-\operatorname{\mathbb{E}}_{k\sim p}B_{k}\right)$ $(\forall k).$ (7)
Equation 6 gives the first-order conditions for the aggregate vote totals $A_{1},\dots,A_{m}$ in pure-strategy equilibrium when the total welfare of each alternative is given by $W_{1},\dots,W_{m}$. Equation 7 gives the best responses $a_{1}^{\text{mech}},\dots,a_{m}^{\text{mech}}$ to a particular aggregate total $A_{1},\dots,A_{k}$ by an agent with value vector $\mathbf{B}$. The equations generally form a fixed-point problem.
Synthetic Players QTM.
The Synthetic Players QTM proceeds as follows.
@@ 3.1 Model and Mechanism
    Set $c\geq\tfrac{1}{2}\max_{i,k}v_{k}^{i}$.
@@
    Collect votes $\mathbf{a}^{1},\dots,\mathbf{a}^{n}$ from the agents.
@@
    Collect payments based on votes and redistribute as in the QTM, i.e. Equation 2.
@@
    Define $A_{k}=a_{k}^{\text{mech}}+\sum_{i}a_{k}^{i}$ and solve Equations 6 and 7 for $a_{1}^{\text{mech}},\dots,a_{m}^{\text{mech}}$ and $A_{1},\dots,A_{m}$.
  5. 5.
    Draw an alternative from $\mathbf{p}$ defined by $p_{k}=e^{A_{k}}/\sum_{\ell}e^{A_{\ell}}$.
The idea of the mechanism is to simulate a set of synthetic players whose total values are $\mathbf{B}$ and total votes are $\mathbf{a}^{\text{mech}}$. (The payments of the synthetic players are irrelevant and can be set to zero, as shown in our analysis of the second variant below.) In an equilibrium of the original QTM, given the votes of the real agents, the synthetic agents’ best responses and the QTM output are given by Equations 6 and 7.
Difficulty of analysis.
The catch is that the Practical Synthetic Players QTM gives the (real) agents commitment power. In an equilibrium of the original QTM, if a real agent deviates to another strategy, the other agents would continue to play the equilibrium. But in this mechanism, the synthetic agents calculate their strategies in response to the strategies of the real agents. This opens the possibility for manipulation. To see this, note that in the two-alternative case, a succinct description of the mechanism is to collect $\mathbf{a}$, then solve the fixed-point problem
$p_{1}=\frac{e^{\left(\sum_{i}a_{1}^{i}\right)+\tfrac{p_{1}p_{2}}{2c}(B_{1}-B_{2})}}{e^{\left(\sum_{i}a_{1}^{i}\right)+\tfrac{p_{1}p_{2}}{2c}(B_{1}-B_{2})}+e^{\left(\sum_{i}a_{2}^{i}\right)+\tfrac{p_{1}p_{2}}{2c}(B_{2}-B_{1})}},$ (8)
with $p_{2}=1-p_{1}$. We conjecture this mechanism has roughly the same Price of Anarchy guarantees as the impractical variant below. However, the change has an undeniable strategic impact. A tempting modification would be to replace each $p_{1}p_{2}$ in (8) with an estimated constant, such as $\tfrac{1}{4}$, removing the ability of agents to influence the impact of the external welfare terms and replacing the fixed-point problem with a simple equation. However, for that modification, there exist examples with low-welfare equilibria in which the influence of $\mathbf{B}$ dominates the decision, although $\mathbf{V}$ has more importance.
Synthetic Players QTM – Impractical Variant.
We now propose a variant that is less practical, but has strong theoretical guarantees. Specifically, we assume the mechanism has knowledge of $\mathbf{V}=(V_{1},\dots,V_{m})$, the total values of the agents. (The mechanism does not require knowledge of individual values beyond an upper bound on the maximum, as in the QTM.) With this, we flip the commitment power around: the mechanism first commits to the votes of the synthetic agents, then the real agents best-respond.
This variant may be practical in some settings where the mechanism does have approximate knowledge of $\mathbf{V}$. For example, the decisionmaking process may include discussions, straw polls, etc. However, in general, we view the assumption as being an impractical one made for the sake of analysis. We view our results for the impractical variant as giving strong evidence for the performance of the Practical variant or other similar mechanisms based on QTM.
The Impractical Variant of the Synthetic Players QTM proceeds as follows.
  1. 1.
    Assume knowledge of the value totals $\mathbf{V}$ (though not individual values).
  2. 2.
    Set $c\geq\tfrac{1}{2}\max_{i,k}v_{k}^{i}$.
  3. 3.
    Compute a solution $\mathbf{A},\mathbf{p}$ that satisfies Equation 6.
  4. 4.
    Announce $\mathbf{A}$; calculate and announce $a_{1}^{\text{mech}},\dots,a_{m}^{\text{mech}}$ from Equation 7.
  5. 5.
    Run the QTM, but increase the total votes of each alternative $k$ by $a_{k}^{\text{mech}}$ before taking the softmax. That is, collect votes $\mathbf{a}=\mathbf{a}^{1},\dots,\mathbf{a}^{n}$ from the participants, assign payments based on $\mathbf{a}$ exactly as in the QTM, and select an alternative according to $\mathbf{p}^{\prime}$ with
    $p_{k}^{\prime}=\frac{e^{a_{k}^{\text{mech}}+\sum_{i}a_{k}^{i}}}{\sum_{\ell}e^{a_{\ell}^{\text{mech}}+\sum_{i}a_{\ell}^{i}}}.$
3.2 Analysis
Focusing on the two-alternative case, we show that the Impractical Variant inherits the welfare guarantee of the QTM. We assume for equilibrium selection that if there are multiple pure-strategy equilibria, agents select the “focal” equilibrium with vote totals $\mathbf{A}$ announced by the mechanism. Denote $\operatorname{\mathbb{E}}_{k\sim p}W_{k}=\sum_{k=1}^{m}p_{k}W_{k}$ as the expected total welfare of the mechanism. In this case, our Price of Anarchy results for the QTM can be transferred to this setting.
Proposition 4.
[…]

### v2

@@ 3 Quadratic Transfers Mechanism
In this section, we give results, including a Price of Anarchy bound, for the Quadratic Transfers Mechanism (QTM) for the public projects problem. These results complement existing analyses (Eguia et al., 2019, 2023), and will also serve as a foundation for our results in later sections combining predictions and preferences.
@@ 3.1 Model
In this section, we consider the special case of our setting without external welfare; that is, the classic public projects setting. There are $n$ agents with preference profile $\mathbf{v}^{i}$ who wish to select one of $m\geq 2$ alternatives. In this section, the welfare of alternative $k$ is $V_{k}=\sum_{i=1}^{n}v_{k}^{i}$. Without loss of generality, in this section we number the alternatives such that $V_{1}\geq V_{2}\geq\ldots\geq V_{m}$. This is a special case of the setting in Section 2, where it is known that $B_{k}=0$ for all alternatives $k$ and agents only participate in the second stage.
A decisionmaking mechanism takes the actions of agents as input and outputs a probability distribution $\mathbf{p}\in\Delta_{m}$, along with a net payment from each agent. We assume quasilinear utility: if the mechanism’s output is $\mathbf{p}$ and if agent $i$ makes a net payment $\pi^{i}$, then $i$’s expected utility is given by $\sum_{k}p_{k}v_{k}^{i}-\pi^{i}$. The social welfare of the mechanism is $\sum_{k}p_{k}V_{k}$, the expected aggregate value produced. The optimal or “first-best” social welfare is $V_{1}$, the aggregate value of the best alternative.
@@
In the Quadratic Transfers Mechanism (QTM) (Eguia et al., 2019, 2023), each agent $i$ simultaneously and privately submits a vector $\mathbf{a}^{i}\in\operatorname{\mathbb{R}}^{m}$, where the $k$th entry $a_{k}^{i}$ represents a number of “votes” for alternative $k$. Votes are allowed to be negative, representing voting against that alternative. We let $A_{k}:=\sum_{i}a_{k}^{i}$ denote the aggregate votes for alternative $k$.
@@
3.2 Price of Anarchy: Two Alternatives
In this section, we present our results for the QTM on $m=2$ alternatives. Conceptually similar results have already been shown by Eguia et al. (2019, 2023) (see Section 1.2), but they focus more on the challenges of the $m\geq 3$ setting and have different distributional and model assumptions than we do here. We note that this section relies on an analysis of equilibria provided in Appendix A.1. Specifically, we first show pure Nash equilibria exist given a lower bound on $c$, and prove that for $m=2$ alternatives equilibria are unique:
@@ 3.2 Price of Anarchy: Two Alternatives
In the QTM on agent value profile $\mathbf{v}$, if the mechanism chooses $c\geq\tfrac{1}{2}\max_{i,k}v_{k}^{i}$, then a pure-strategy Nash equilibrium exists. If $m=2$, it is unique.
The result is proven as Propositions 4 and 5 in Appendix A.
Then we observe several properties of the first order conditions for agent votes. We also characterize the revenue (before redistributions) in terms of “disagreement” in the type profile in Appendix A.2, and briefly touch on the case where there are more than two alternatives in Appendix A.3.
Price of Anarchy result.
Recall that $V_{1}$ is the aggregate value of the better alternative, i.e. optimal welfare, and $V_{2}$ is the aggregate value of the other. Our main result, Theorem 1, is that the QTM’s pure-strategy Price of Anarchy (pPoA) is always at least $\tfrac{1}{2}$ and approaches $1$ rapidly as the spread of the game — the ratio of $V_{1}$ to the largest individual value — grows. We also show in Corollary 1 a faster rate of convergence to $1$ when the gap — the difference between $V_{1}$ and $V_{2}$ — grows. Finally, we will characterize the total payments of the QTM in terms of the amount of “disagreement” of the participants.
@@
$\text{pPoA}(\mathbf{v})\geq\max\left\{\frac{1}{2}~{},~{}~{}1-\left(\frac{2}{T}\right)^{2/5}\right\}.$
@@
$\text{pPoA}(\mathbf{v})\geq p_{1}\geq\max\left\{\frac{1}{2}~{},~{}~{}1-\left(\frac{4}{G}\right)^{2/3}\right\}.$
The case of $m\geq 3$ is highly complex. Eguia et al. (2019, 2023) give an asymptotic analysis for a Bayesian setting showing a Price of Anarchy approaching one. For a Nash equilibrium setting with no assumptions, we do not obtain such a result, but we do have a Price of Anarchy bound that does not depend at all on the number of participants or their values.
Proposition 2.
For the QTM on $m$ alternatives, for any choice of $c\geq\tfrac{1}{2}\max_{i,k}v_{k}^{i}$, we have $\text{pPoA}\geq\frac{1}{m}$.
The proof appears in Appendix A.3. It involves iteratively dividing the alternatives into a group with “high” welfare (higher than the mechanism’s expectation) and the remainder, an idea from Eguia et al. (2023) used in a different way.
4 Incorporating External Welfare
In this section, we propose mechanisms for public projects with external welfare, meaning estimates or projections of an externality of the project on non-decisionmakers. We will suppose that the external welfare impacts are fixed and common knowledge of the agents and the mechanism. Thus we are in a special case of the setting in Section 2 without an aggregation stage, where the external welfare impacts are fully known in advance, i.e. $B_{k}=\hat{B}_{k}=B_{k}^{*}$ is the external welfare impact of alternative $k$. In Section 5, we will extend to the general case where external welfare impacts are forecasts from (manipulable) aggregation mechanisms.
Model.
We define the problem of Public Projects with External Welfare as an extension of the standard public projects setting where selecting alternative $k$ causes a known, fixed external welfare impact $B_{k}$. The $n$ agents continue to have values $\{v_{k}^{i}\}$, and the total value of alternative $k$ continues to be $V_{k}=\sum_{i=1}^{n}v_{k}^{i}$. The welfare if $k$ is chosen is $W_{k}\coloneqq V_{k}+B_{k}$, and we assume without loss of generality that alternatives are numbered according to $W_{1}\geq W_{2}\geq\ldots\geq W_{m}$. We let $\text{pPoA}(\mathbf{v},\mathbf{B})$ be the pure-strategy Price of Anarchy of our mechanism in the setting $(\mathbf{v},\mathbf{B})$, i.e. the worst-case ratio of welfare in pure-strategy Nash equilibrium to $W_{1}$.
Mechanism.
We propose the Synthetic Players Quadratic Transfer Mechanism for public projects with external welfare. The mechanism works by creating a set of synthetic players whose preferences reflect the external welfare, then playing on their behalf in the QTM alongside the real participants. It proceeds as follows:
  1. 1.
    The mechanism is given $B_{1},\dots,B_{m}$.
  2. 2.
    Set $c\geq\tfrac{1}{2}\max_{i\leq n,k}v_{k}^{i}$.
  3. 3.
    Define $\hat{n}\geq\frac{\max_{k}B_{k}}{2c}$ “synthetic players,” each with a value profile of $\frac{1}{\hat{n}}(B_{1},\dots,B_{m})$.
  4. 4.
    Run the QTM on real agents $1,\dots,n$ and synthetic agents $\hat{1},\dots,\hat{n}$, where the mechanism plays on behalf of the synthetic players. Do not redistribute the payments of the synthetic players.
We obtain the following Price of Anarchy guarantee for the mechanism:
Proposition 3.
Suppose the QTM with $m$ alternatives on profile $\mathbf{v}$ guarantee $\text{pPoA}{(\mathbf{v})}\geq C(\mathbf{v}).$ Then the Synthetic Players QTM on $m$ alternatives, for all external welfare vectors $\mathbf{B}$, achieves budget balance and guarantees $\text{pPoA}{(\mathbf{v},\mathbf{B})}\geq C(\mathbf{v}).$
Proof.
Budget balance follows immediately from budget balance of the QTM. For each synthetic player $j=\hat{1},\dots,\hat{n}$, let her value vector be $\mathbf{\hat{v}}^{j}=\tfrac{1}{\hat{n}}\mathbf{B}$. Observe $c\geq\tfrac{1}{2}\max_{j,k}\hat{v}_{k}^{j}$, and by construction $c\geq\tfrac{1}{2}\max_{i\leq n,k}v_{k}^{i}$, so that a pure-strategy Nash equilibrium is guaranteed to exist by Proposition 1. Moreover, the welfare of each alternative $k$ in the game is $W_{k}$. The result follows. ∎
An immediate corollary, using the Price of Anarchy results in Theorem 1 and Corollary 1 for $m=2$, is:
Corollary 2.
For the Synthetic Players QTM on two alternatives, for any profile $\mathbf{v}$, there exists a choice of parameter $c$ such that, for all external welfare vectors $\mathbf{B}$, the mechanism achieves budget balance and
  1. 1.
    $\text{pPoA}(\mathbf{v},\mathbf{B})\geq\max\{\tfrac{1}{2},1-(\tfrac{2}{T})^{2/5}\}$, where $T=\frac{W_{1}}{\max_{i,k}v_{k}^{i}}$ is the spread;
  2. 2.
    $\text{pPoA}(\mathbf{v},\mathbf{B})\geq p_{1}\geq\max\{\tfrac{1}{2},1-\left(\tfrac{4}{G}\right)^{2/3}\}$, where $G=\frac{W_{1}-W_{2}}{max_{i,k}v_{k}^{i}}$ is the gap. [^3] Note that the inequality in terms of $p_{1}$ holds since any guarantees of the QTM for $\mathbf{p}$ extend to the external welfare setting by the proof of Proposition 3.
Practical concerns and mechanism variants.
To implement the mechanism, one needs to compute an equilibrium strategy for the synthetic players. Although this only requires knowledge of the vote totals of the other players – not their individual strategies – it may be impractical in many cases. One mitigating factor is that, in the two-alternative case, equilibrium is unique (see Proposition 1), so equilibrium selection is not a concern.
We also suggest two potential variants of the mechanism that remove the requirement for the mechanism to compute an equilibrium. The first is a sequential voting game similar to the sequential first-price public projects mechanism of Lucier et al. (2013). We ask the “real” participants to cast votes publicly one at a time, with the mechanism going last on behalf of the synthetic players. The second is a two-stage version, in which all “real” participants cast their votes in stage one simultaneously, and then the mechanism adds the synthetic votes in stage two. In both of these suggestions, the mechanism computes the synthetic players’ total votes and the output of the mechanism using the first-order conditions. In the two-alternative case, for example, given the real players’ votes $\{a_{k}^{i}\}$, the mechanism solves the fixed-point problem
$p_{1}=\frac{e^{\left(\sum_{i}a_{1}^{i}\right)+\tfrac{p_{1}p_{2}}{2c}(B_{1}-B_{2})}}{e^{\left(\sum_{i}a_{1}^{i}\right)+\tfrac{p_{1}p_{2}}{2c}(B_{1}-B_{2})}+e^{\left(\sum_{i}a_{2}^{i}\right)+\tfrac{p_{1}p_{2}}{2c}(B_{2}-B_{1})}}.$
In practice, we expect both variants to be practical. In particular, in large games, we do not expect participants to be able to strategize significantly to influence the result of the mechanism. However, strategic analysis of these variants is left to future work.
5 Incorporating Predictions
Armed with the analyses in Sections 3 and 4, we now return to the setting of public projects with predictions. As a reminder from Section 2, an aggregation mechanism is operated first in order to advise a decisionmaking mechanism, which operates second. The first stage produces as output forecasts $\mathbf{\hat{B}}=(\hat{B}_{1},\dots,\hat{B}_{m})$. The second stage is to run the QTM with external welfare impacts $\mathbf{\hat{B}}$, as defined in Section 4, which outputs a probability distribution $\mathbf{p}$ over the alternatives and selects from that distribution.
5.1 The Two-Stage Game
We begin by explicitly defining strategies and utilities in the two-stage game. We remind the reader that the external welfare impact conditional on choosing alternative $k$, $B_{k}$, is now a random variable. Agent $i$ receives a signal $S_{i}$ drawn jointly with $(B_{1},\dots,B_{m})$ from a common-knowledge prior $\mathcal{D}$. When an agent acts in the first stage, we refer to them as a forecaster; in the second stage, as a decisionmaker.
As a forecaster, agent $i$ submits a prediction vector $\mathbf{\hat{b}}^{i}$ according to her strategy $\sigma^{i,1}:S_{i}\to\operatorname{\mathbb{R}}^{m}$. Let $\mathbf{\hat{b}}=(\mathbf{\hat{b}}^{1},\ldots,\mathbf{\hat{b}}^{n})$ be the set of prediction vectors across forecasters. She is paid according to some function $\pi^{i,1}(\mathbf{\hat{b}}^{i},\mathbf{p},k,b_{k})$ depending on the eventual observation of the actual outcome $B_{k}$ of the alternative chosen by the joint mechanism. Thus agent $i$’s expected utility $u^{i,1}$ in this stage corresponds to her expected payment: $u^{i,1}(\mathbf{\hat{b}}^{i},\mathbf{p})=\sum_{k}p_{k}\operatorname{\mathbb{E}}_{b_{k}\sim\mathcal{D}_{k}}\pi^{i,1}(\sigma^{i,1},\mathbf{p},k,b_{k})$. As a decisionmaker, agent $i$ also submits a vote vector $\mathbf{a}^{i}$ to the decisionmaking mechanism according to strategy $\sigma^{i,2}:(S_{i},{\mathbf{\hat{b}}},\mathbf{\hat{B}},\mathbf{a}^{-i})\to\operatorname{\mathbb{R}}^{m}$. She receives some payment $\pi^{i,2}(\mathbf{a})$, so her expected utility $u^{i,2}$ in this stage (under the assumption of quasilinear utilities) is $u^{i,2}(\mathbf{a},\mathbf{p})=\sum_{k}p_{k}v_{k}+\pi_{2}^{i}(\mathbf{a}).$
@@ 5.1 The Two-Stage Game
Mechanism overview.
The specific mechanism implementation we propose, the Synthetic players QUAdratic transfer mechanism with Predictions (SQUAP), is as follows:
@@
    Forecasters participate in the aggregation mechanism, which produces $\mathbf{\hat{B}}=(\hat{B}_{1},\dots,\hat{B}_{m})$.
@@
    We run the Synthetic Players QTM defined in Section 4 using external welfare impacts $\mathbf{\hat{B}}$.
@@
    The output distribution is $\vec{p}$ and the selected alternative is $k$.
@@
    Later, $B_{k}$ is observed and the aggregation mechanism rewards participants based on $\vec{p},k,B_{k}$.
Redistribution of QTM payments has the potential, at least in theory, to cause incentive problems. For theoretical analysis, we assume that the mechanism keeps the payments or “burns” them. In practice, designers may choose to redistribute them or use them to subsidize the aggregation mechanism, as the incentive concern is likely low, but we leave that analysis to future work.
5.2 SQUAP Price of Anarchy
Incentive alignment.
Perhaps surprisingly, we only require two natural properties of an aggregation mechanism in order to prove welfare guarantees in equilibrium of the combined mechanism. These properties address the two possible incentive problems previously mentioned: a voter may manipulate the predictions in order to boost some alternatives’ chances of being chosen, and a forecaster may manipulate the decisionmaking mechanism in order to increase the rewards for their predictions. First is the incentive of forecasters to manipulate the decision in order to gain higher prediction rewards.
Definition 1.
An aggregation mechanism is alternative-independent if the expected reward $u^{i,1}$ of any forecaster $i$ does not depend on the distribution $\mathbf{p}$ from which the final alternative is chosen, as long as $\mathbf{p}$ has full support.
We will see that an importance-weighting technique due to Chen et al. (2011) allows one to achieve alternative-independence, so that prediction rewards are independent of the decisionmaking mechanism (at least in expectation.) Under alternative independence, a forecaster does not gain increased information rewards by changing the decisionmaking mechanism’s output $\mathbf{p}$ because her utility is now only a function of her prediction, i.e. $u^{i,1}(\mathbf{\hat{b}}^{i})$. It follows that an agent’s strategy $\sigma^{i,2}$ in any QTM subgame is to maximize utility in that complete-information subgame exactly as in Section 4, when external welfare was fixed and known.
Next is the incentive of decisionmakers to manipulate their predictions $\mathbf{\hat{b}}$ in order to boost the chances of their preferred alternative(s). In general, agents can gain from manipulating the estimated external welfare even at a cost. However, we will be able to show that the QTM’s welfare guarantees are robust to relatively small manipulations. Therefore, we only need that large manipulations are prohibitively costly. We formalize this costliness as follows. Say that a player is $x$-best-responding in a game if they can improve their net expected payoff $u^{i}$ by at most $x$ by switching to another strategy.
Definition 2.
In an aggregation mechanism, let $\mathbf{B^{*}}$ be the output of the mechanism if all forecasters are truthful and let $\mathbf{\hat{B}}$ be the output under strategic behavior. Given some $x>0$ and $\alpha>0$, we say the aggregation mechanism is $(\alpha,x)$-robust if, for all strategy profiles $\vec{\sigma}$ where all participants are $x$-best-responding,
$\max_{k}|\hat{B}_{k}-B_{k}^{*}|\leq\alpha x.$
Solution concept.
Because SQUAP is a dynamic game, we consider a refinement of Nash equilibria: a strategy profile $\vec{\sigma}$ is in sub-game perfect pure equilibrium if it is in pure Bayes Nash equilibrium and, in addition, the strategies $\vec{\sigma}^{\,\boldsymbol{\cdot},2}$ played at each subgame consisting of the Synthetic Players QTM are in pure Nash equilibrium. In a setting defined by valuations $\mathbf{v}$ and information structure $\mathcal{D}$, we let $\text{pPNE}(\mathbf{v},\mathcal{D})$ denote the set of distributions $\mathbf{p}\in\Delta_{m}$ induced by the mechanism in any sub-game perfect pure equilibrium. Then we define $\text{pPoA}(\mathbf{v},\mathcal{D})$ as the minimum ratio of the welfare in sub-game perfect pure equilibrium to $W_{1}=V_{1}+B_{1}^{*}$, i.e.
$\text{pPoA}{(\mathbf{v},\mathcal{D})}=\frac{\min_{\mathbf{p}\in\text{pPNE}(\mathbf{v},\mathcal{D})}\sum_{k}p_{k}W_{k}}{W_{1}}.$ (4)
PoA results.
To prove our main Price of Anarchy results, we first use the strategic properties of alternative-independence and deviation robustness to show that, in equilibrium, the estimates $\mathbf{\hat{B}}$ are reasonably accurate.
Lemma 1.
In any sub-game perfect pure equilibrium of SQUAP, if the aggregation mechanism is alternative-independent and $(\alpha,x)$-robust, then its output satisfies for all $k$ the bound $|\hat{B}_{k}-B_{k}^{*}|\leq\alpha\max_{i,k}v_{k}^{i}$.